# PlaneRecTR++: Unified Query Learning for Joint 3D Planar Reconstruction and Pose Estimation

### v3

@@ 4.1 Transformer-based Unified Query Learning for Single-view Plane Recovery
Inspired by the recent successes of DETR [20] and Mask2Former [21] in object detection and instance segmentation, we find that, it is not impossible to tackle the challenging monocular planar reconstruction task using a single, compact and unified framework, thanks to the merits of query-based reasoning to enable joint modeling of multiple tasks.
@@
As shown in Figure 3, intra-frame plane query learning component consists of three main modules: (1) A pixel-level module to learn dense pixel-wise deep embedding of the input colour image. (2) A Transformer-based unified query learning module to jointly predict, for each of $N$ learnable plane queries, its corresponding plane embeddings $\mathcal{E}_{\text{plane}}$ as well as four target properties, including plane classification probability $p_{i}$, plane parameter $n_{i}$, mask embedding, and depth embedding ($i\in[1,2,...,N]$). Specifically, $p_{i}$ is the probability to judge whether the $i^{\text{th}}$ query corresponds to a plane or not; $n_{i}\doteq\tilde{n_{i}}/d_{i}\in\mathbb{R}^{3}$, where $\tilde{n_{i}}\in\mathbb{R}^{3}$ is its plane normal and $d_{i}$ is the distance from the $i^{\text{th}}$ plane to camera center, i.e., offset. (3) A plane-level module to generate plane-level mask $m_{i}$ and plane-level depth $d_{i}$ through mask and depth embedding ($i\in[1,2,...,N]$). We then remove non-plane query hypothesis while combining the remaining ones for the final image-wise plane recovery. These three modules will be described in details bellow.
@@
Pixel-Level Module.   Given an image of size $H\times W$ as input, we use the pre-trained ResNet-50 [43] as the backbone model to extract dense image feature maps, unless otherwise mentioned. Similar to Mask2Former, a multi-scale convolutional pixel decoder is used to produce a set of dense feature maps with four scales, denoted as follows:
$\mathbb{F}$ =$=\{F_{1}\in\mathbb{R}^{C_{1}\times H/32\times W/32},F_{2}\in\mathbb{R}^{C_{2}\times H/16\times W/16},$ (1)
$F_{3}\in\mathbb{R}^{C_{3}\times H/8\times W/8},\mathcal{E}_{\text{pixel}}\in\mathbb{R}^{C_{\mathcal{E}}\times H_{\mathcal{E}}\times W_{\mathcal{E}}}\}.$
@@
Among these, the first three feature maps $\{F_{1},F_{2},F_{3}\}$ are fed to the Transformer module, while the last one $\mathcal{E}_{\text{pixel}}$, a dense per-pixel embedding of resolution $H_{\mathcal{E}}=H/4$ and $W_{\mathcal{E}}=W/4$, is exclusively used for computing plane-level binary masks and plane-level depths.
@@
Transformer Module.   We use the Transformer decoder with masked attention proposed in [21], which computes unified plane embeddings $\mathcal{E}_{\text{plane}}\in\mathbb{R}^{N\times C_{\mathcal{E}}}$ from above mentioned multi-scale feature maps $\{F_{1},F_{2},F_{3}\}$ and $N$ learnable plane queries. The predicted $\mathcal{E}_{\text{plane}}$ are then independently projected to four target properties by four different multi-layer perceptrons (MLPs). Overall, the Transformer module predicts required planar attributes through $N$ plane queries (upper left part in Figure 3).
@@
Plane-Level Module.   As shown in the upper right part of Figure 3, we obtain a dense plane-level binary mask $m_{i}\in[0,1]^{H_{\mathcal{E}}\times W_{\mathcal{E}}}$/depth prediction $d_{i}\in\mathbb{R}^{H_{\mathcal{E}}\times W_{\mathcal{E}}}$ by a dot product between the $i^{\text{th}}$ mask/depth embedding and the dense per-pixel embedding $\mathcal{E}_{\text{pixel}}$ from previous two modules, respectively.
Please note that we have also investigated the idea of learning two individual pixel decoders respectively for semantic and depth dense embeddings, in order to separately predict planar masks and depths. However, we did not observe a clear performance improvement, and therefore stick to current set-up with a shared $\mathcal{E}_{\text{pixel}}$ for the sake of efficiency and simplicity.
We finally obtain $N$ plane-level predictions $\{y_{i}=(p_{i},n_{i},m_{i},d_{i})\}_{i=1}^{N}$ by the plane-level module, and each plane-level prediction contains all the necessary information to recover a possible 3D plane.
@@ 4.2 Training Objective and Configuration
Plane-level Depth Training.   Previous methods tend to predict a global image-wise depth by separate network branches to calculate plane offset [13] or use depth as additional cues to formulate segmentation masks [14, 15]. In contrast, our method tries to achieve mutual benefits between planar semantic and geometric reasoning, we leverage joint query learning to unify all components of plane recovery in a concise multi-task manner. As a result, we explicitly predict dense plane-level depths, binary-masks, plane probabilities and parameters from a shared feature space, which is produced and refined via attention mechanism of the Transformer.
@@
[…]
@@
Loss Functions.   After bipartite matching, the final training objectives $L$ is composed of following four parts:
$\mathcal{L}=\sum\limits_{i=1}^{M}\left(\lambda\mathcal{L}_{\text{cls }}^{(i)}+\mathcal{L}_{\text{param }}^{(i)}+\mathcal{L}_{\text{mask}}^{(i)}+\lambda\mathcal{L}_{\text{depth}}^{(i)}\right),$ (4)
@@
However, different from PlaneTR [15], the left two loss terms in our paper $\mathcal{L}_{\text{mask}}$ and $\mathcal{L}_{\text{depth}}$ are designed to explicitly learn dense planar masks and depths. Specifically, we introduce the plane segmentation mask prediction loss, as a combination of a cross-entropy loss and a dice loss:
@@
where $\beta_{1}=5$ and $\beta_{2}=5$. The depth prediction loss is in a typical $L_{1}$ form, penalizing the discrepancy of depth value within planar regions:
$\mathcal{L}_{\text{depth }}^{(i)}=\mathbbm{1}_{\{\hat{p}_{i}=1\}}L_{1}\left(\hat{{d}}_{i},{d}_{\sigma(i)}\hat{m}_{i}\right).$ (5)
@@ 4.3 Inference Process of Monocular 3D Plane Recovery
In this section, we briefly discuss how to recover 3D planes from our plane-level predictions. For the $N$ plane-level predictions $\{y_{i}\}_{i=1}^{N}$ predicted by the network, we first drop non-plane candidates according to the plane classification probability $p_{i}$, leading to a valid subset of $K$ planes (i.e., $K\leq N$). For each pixel within the planar regions, we calculate the most likely plane index $\mathop{\arg\max}\limits_{i}\{m_{i}\}_{i=1}^{K}$ to obtain its mask id and thus obtain the final image-wise segmentation mask.
In this simple and efficient manner, we can use the network predictions to create 3D plane reconstruction of the input frame. Note that plane-level depths are not involved during inference and we use plane parameters and segmentation to infer planar depths. We experimentally found that this design also leads to more structural and smooth geometric predictions than that relying on direct depth predictions.
Figure 4: Overview of inter-frame plane query learning.
@@ 5 Inter-Frame Plane Query Learning
In this section, we presented how we extend our single-view plane recovery framework PlaneRecTR to a novel multi-view setup, while still retaining the virtue of query-based learning. We introduce an inter-frame query learning component on top of unified plane embeddings of per-frame. A plane aware cross attention layer is proposed to achieve inter-frame plane interactions, within which dual softmax [39, 25] and bilinear attention [45, 25] mechanisms are used to align the intermediate attention structure with a 2D plane correspondence matrix and enable plane-level feature fusion between views. Most importantly, we further modify the key, query, and value forms of standard multi-head attention [16] to effectively utilize the complete representation of unified plane embeddings, which better accommodates multi-view plane properties without requiring any additional inputs such as position encoding [39, 25]. This simple adjustment guarantees that our attention structure truly accomplishes plane matching, allowing the network to spontaneously focus on genuine paired planes from two views.
As illustrated in Figure 4, we employ two plane-aware cross attention layers and an MLP head to construct a simple and lightweight pose regression module. A plane-aware attention layer first takes two-view plane embeddings $\mathcal{E}_{\text{plane}}^{1}$ and $\mathcal{E}_{\text{plane}}^{2}$ as input, and actively learn their correspondences within the network. Subsequently, our model directly regresses a relative camera pose $T=\{R,t\}$ from probabilistic paired plane embeddings. Conceptually, this entire process aligns with the logical framework of solutions based on classical two-view geometry [32], thereby offering enhanced interpretability.
@@ 5.1 Cross Attention for Unified Plane Embeddings
Preliminaries: Standard Cross Attention.   The Transformer cross attention layer [16] updates the input value term by mapping query from the $i$-th input and key-value pair from another $j$-th input through a weighted summation, typically using a scaled dot-product similarity function $\rm{S}(\cdot,\cdot)$. In addition, attention often employs a multi-head strategy [16] where queries, keys, and values (denoted $Q_{i}$, $K_{j}$, $V_{j}$ $\in\mathbb{R}^{N\times C_{\mathcal{E}}}$, respectively) are divided, along channel dimensions, into $N_{h}$ segments $\{q_{i}^{h}\}_{h=1}^{N_{h}}$, $\{k_{j}^{h}\}_{h=1}^{N_{h}}$, $\{v_{j}^{h}\}_{h=1}^{N_{h}}\in\mathbb{R}^{N\times\frac{C_{\mathcal{E}}}{N_{h}}}$, in order to enhance the expressiveness and diversity of results without incurring extra computational costs.
The multi-head similarity function ($\rm{S}$) and cross attention layer ($\rm{MCA}$) are defined in Equations 6 and 7:
$\rm{S}(q_{i}^{h},k_{j}^{h})=\rm{softmax}\left(\frac{q_{i}^{h}{k_{j}^{h}}^{T}}{\sqrt{C_{\mathcal{E}}/N_{h}}},1\right),$ (6)
$\rm{MCA}(Q_{i},K_{j},V_{j})=\rm{Linear}(\rm{Concat}(\{\rm{S}(q_{i}^{h},k_{j}^{h})v_{j}^{h}\}_{h=1}^{N_{h}})),$ (7)
where $\rm{softmax}(\cdot,k)$ applies softmax operation across the $k$-th axis; $\rm{Linear}$ and $\rm{Concat}$ mean linear projection and channel-wise concatenation, respectively.
Our proposed model targets equations 6 and 7 for intuitive and efficient plane-specific modifications, achieving implicit plane matching and direct regression of precise pose.
Plane Correspondence Probability Function.   To overcome the limitations of multi-stage two-view plane reconstruction paradigm [7, 8, 9], here we specifically devised an inter-frame correspondence attention structure to learn reliable plane embeddings, which enable the network to autonomously acquire probabilities of plane correspondence and conduct pose inference in a single forward pass. This design also eliminates the dependency on either ground truth correspondence supervision or initial pose.
@@
$\rm{C}(Q_{i},K_{j})=\rm{softmax}(\frac{Q_{i}K_{j}^{T}}{\sqrt{C_{\mathcal{E}}}},1)\odot\rm{softmax}(\frac{Q_{i}K_{j}^{T}}{\sqrt{C_{\mathcal{E}}}},2).$ (8)
We compute a 2D correspondence matrix $\rm{C}(Q_{i},K_{j})$, where the element at the $m^{th}$ row and $n^{th}$ column $\rm{C}_{mn}(Q_{i},K_{j})$ denotes the probability that the $m^{th}$ plane embedding from the $i^{th}$ image $I_{i}$ corresponds to the $n^{th}$ plane embedding from $j^{th}$ image $I_{j}$, indicating their likelihood of representing the same plane instance. For the task of two-view plane reconstruction, we stick to configurations of $\{i=1,j=2\}$ and $\{i=2,j=1\}$.
@@
Inter-frame Plane Aware Cross Attention.   The standard cross attention offers an efficient approach to selectively utilize one of input plane embedding sequences, but overlooks the interaction between two inputs. Therefore, we adopt bilinear attention [45, 25] to incorporate planar information from both views through the plane correspondence probability distribution $\rm{C}(Q_{i},K_{j})$:
$\rm{PCA}(Q_{i},K_{j},V_{i},V_{j})=$ $\rm{Linear}(\rm{Concat}($
$\{(v_{i}^{h})^{T}\rm{C}(Q_{i},K_{j})v_{j}^{h}\}_{h=1}^{N_{h}})).$ (9)
Unlike query and key terms, the value continues to be divided into $N_{h}$ segments along the feature dimension, maintaining the advantages of multi-head attention. The value segment $v_{j}^{h}\in\frac{C_{\mathcal{E}}}{N_{h}}$ shares the correspondence attention (Equation 8) from unsplit query and key, allowing our model to selectively attend to actual corresponding plane embedding pairs across distinct sub-spaces.
As shown in Figure 4, two parallel plane aware cross attention (Equation 5.1) are parallelly employed to capture integrated features of the corresponding plane from $I_{1}$ to $I_{2}$ as well as from $I_{2}$ to $I_{1}$. It is worth note that pose ViT [25] employs the same features on both sides of bilinear attention matrix. However, we intuitively choose to cross-place embedding sequences of distinct images to model correspondences, thereby enhancing learning efficiency and yielding improved results. We will validate various design disparities through subsequent ablation studies in Section 7.6.

### v4

@@ 4.1 Transformer-based Unified Query Learning for Single-view Plane Recovery
Inspired by the successes of DETR [20] and Mask2Former [21] in object detection and segmentation, we find that, it is feasible to tackle the challenging monocular planar reconstruction task using a single, compact and unified framework, thanks to the merits of query-based reasoning for enabling joint modeling of multiple tasks.
@@
As shown in Figure 3, intra-frame plane query learning component consists of three main modules: (1) A pixel-level module to learn dense pixel-wise deep embedding of the input colour image. (2) A Transformer-based unified query learning module to jointly predict, for each of $N$ learnable plane queries, its corresponding plane embeddings $\mathcal{E}_{\text{plane}}$ as well as four target properties, including plane classification probability $p_{i}$, plane parameter $n_{i}$, mask embedding, and depth embedding ($i\in[1,2,...,N]$). Specifically, $p_{i}$ is the probability to judge whether the $i^{\text{th}}$ query corresponds to a plane or not; $n_{i}\doteq\tilde{n_{i}}/d_{i}\in\mathbb{R}^{3}$, where $\tilde{n_{i}}\in\mathbb{R}^{3}$ is its plane normal and $d_{i}$ is the distance from the $i^{\text{th}}$ plane to camera center, i.e., offset. (3) A plane-level module to generate plane-level mask $m_{i}$ and plane-level depth $d_{i}$ through mask and depth embedding ($i\in[1,2,...,N]$). We then remove non-plane query hypothesis while combining the remaining ones for the final image-wise plane recovery. These three modules will be described in detail below.
@@
Pixel-Level Module.  Given an input image of size $H\times W$, we use the pre-trained ResNet-50 [43] as backbone to extract dense image feature maps, unless otherwise mentioned. Subsequently, a multi-scale convolutional pixel decoder [21] is used to produce a set of dense feature maps with four scales, denoted as follows:
$\mathbb{F}$ $=\{F_{1}\in\mathbb{R}^{C_{1}\times H/32\times W/32},F_{2}\in\mathbb{R}^{C_{2}\times H/16\times W/16},$ (1)
$F_{3}\in\mathbb{R}^{C_{3}\times H/8\times W/8},\mathcal{E}_{\text{pixel}}\in\mathbb{R}^{C_{\mathcal{E}}\times H_{\mathcal{E}}\times W_{\mathcal{E}}}\},$
@@
where $C_{1}$, $C_{2}$, $_{3}, CC_{\mathcal{E}}$ are feature dimensions. The first three feature maps $\{F_{1},F_{2},F_{3}\}$ are fed to the Transformer module, while the last one $\mathcal{E}_{\text{pixel}}$, a dense per-pixel embedding of resolution $H_{\mathcal{E}}=H/4$ and $W_{\mathcal{E}}=W/4$, is exclusively used for computing plane-level binary masks and plane-level depths.
@@
Transformer Module.  We use the Transformer decoder with masked attention proposed in [21], which computes unified plane embeddings $\mathcal{E}_{\text{plane}}\in\mathbb{R}^{N\times C_{\mathcal{E}}}$ from above mentioned multi-scale feature maps $\{F_{1},F_{2},F_{3}\}$ and $N$ learnable plane queries. The predicted $\mathcal{E}_{\text{plane}}$ are then independently projected to four target properties by four MLPs. Overall, the Transformer module predicts required planar attributes through $N$ plane queries.
@@
Plane-Level Module.  As shown in Figure 3, we obtain a dense plane-level binary mask $m_{i}\in[0,1]^{H_{\mathcal{E}}\times W_{\mathcal{E}}}$/depth prediction $d_{i}\in\mathbb{R}^{H_{\mathcal{E}}\times W_{\mathcal{E}}}$ by a dot product between the $i^{\text{th}}$ mask/depth embedding and the dense per-pixel embedding $\mathcal{E}_{\text{pixel}}$ from previous two modules, respectively. We finally obtain $N$ plane-level predictions $\{y_{i}=(p_{i},n_{i},m_{i},d_{i})\}_{i=1}^{N}$, each of which contains all the necessary information to recover a possible 3D plane.
@@ 4.2 Training Objective and Configuration
Plane-level Depth Training.  Previous methods tend to predict a global image-wise depth by separate network branches to calculate plane offset [13] or use depth as additional cues to formulate segmentation [14, 15]. In contrast, our method tries to achieve mutual benefits between planar semantic and geometric reasoning, and leverages learnable plane queries to unify all components of plane recovery in a concise multi-task manner. As a result, we explicitly predict dense plane-level depths, binary-masks, plane probabilities and parameters from a shared feature space, which is produced and refined via attention mechanism of the Transformer.
@@
$\hat{\sigma}=\underset{\sigma}{\arg\min}\sum_{i=1}^{N}D\left(\hat{y}_{i},y_{\sigma(i)}\right),$ (2)
@@
Loss Functions.  After bipartite matching, the final training objective $L$ is composed of the following four parts:
$\small\mathcal{L}=\sum\limits_{i=1}^{M}\left(\lambda\mathcal{L}_{\text{cls }}^{(i)}+\mathcal{L}_{\text{param }}^{(i)}+\mathcal{L}_{\text{mask}}^{(i)}+\lambda\mathcal{L}_{\text{depth}}^{(i)}\right),$ (4)
@@
However, different from PlaneTR [15], the remaining two loss terms in our paper $\mathcal{L}_{\text{mask}}$ and $\mathcal{L}_{\text{depth}}$ are designed to explicitly learn dense planar masks and depths. Specifically, we introduce the plane segmentation mask prediction loss, as a combination of a cross-entropy loss and a dice loss:
@@
where $\beta_{1},\beta_{2}=5$. The depth loss is in a typical $L_{1}$ form, penalizing the discrepancy of depth value within planar regions:
$\mathcal{L}_{\text{depth }}^{(i)}=\mathbbm{1}_{\{\hat{p}_{i}=1\}}L_{1}\left(\hat{{d}}_{i},{d}_{\sigma(i)}\hat{m}_{i}\right),$ (5)
@@ 4.3 Inference Process of Monocular 3D Plane Recovery
For the $N$ plane-level predictions $\{y_{i}\}_{i=1}^{N}$ predicted by the network, we first drop non-plane candidates according to the plane classification probability $p_{i}$, leading to a valid subset of $K$ planes (i.e., $K\leq N$). For each pixel within planar regions, we calculate the most likely plane index $\mathop{\arg\max}\limits_{i}\{m_{i}\}_{i=1}^{K}$ to obtain the final image-wise segmentation mask. Note that plane-level depths are not involved during inference and we use plane parameters and segmentation to infer planar depths. We experimentally found that this design also leads to more structural and smooth geometric predictions than that relying on direct depth predictions.
Figure 4: Overview of inter-frame plane query learning. Our inter-frame component encompasses two parts: (1) Two parallel plane-aware cross attention layers to predict potential cross-view plane correspondences without any association labels; (2) A simple MLP-based pose regressor to directly estimate the relative camera pose.
@@ 5 Inter-Frame Plane Query Learning
In this section, we present how we extend our monocular framework PlaneRecTR to a novel multi-view setup, while still retaining the virtue of query-based learning. We introduce an inter-frame query learning component on top of unified plane embeddings of per-frame. A plane aware cross attention layer is proposed to achieve inter-frame plane interactions, within which dual softmax [39, 25] and bilinear attention [45, 25] mechanisms are used to align the intermediate attention structure with a plane correspondence matrix and enable plane-level feature fusion between views. Most importantly, we further modify the key, query, and value forms of standard multi-head attention [16] to effectively utilize the complete representation of unified plane embeddings, which better accommodates multi-view plane properties without requiring any additional inputs such as position encoding [39, 25]. This simple adjustment guarantees that our attention structure truly accomplishes plane matching, allowing the network to spontaneously focus on genuine paired planes from two views.
As illustrated in Figure 4, we employ two plane-aware cross attention layers and an MLP head to construct a simple and lightweight pose regression module. A plane-aware attention layer first takes two-view plane embeddings $\mathcal{E}_{\text{plane}}^{1}$ and $\mathcal{E}_{\text{plane}}^{2}$ as input, and actively learns their correspondences within the network. Subsequently, our model directly regresses a relative camera pose from probabilistic paired plane embeddings. Conceptually, this entire process aligns with the logical framework of solutions based on classical two-view geometry [32], thereby offering enhanced interpretability.
@@ 5.1 Cross Attention for Unified Plane Embeddings
Preliminaries: Standard Cross Attention.  The Transformer cross attention layer [16] updates the input value term by mapping query from the $i$-th input and key-value pair from another $j$-th input through a weighted summation, typically using a scaled dot-product similarity function $\operatorname{S}(\cdot,\cdot)$. In addition, attention often employs a multi-head strategy [16] where query, key, and value (denoted $Q_{i}$, $K_{j}$, $V_{j}$ $\in\mathbb{R}^{N\times C_{\mathcal{E}}}$, respectively) are divided, along channel dimensions, into $N_{h}$ segments $\{q_{i}^{h}\}_{h=1}^{N_{h}}$, $\{k_{j}^{h}\}_{h=1}^{N_{h}}$, $\{v_{j}^{h}\}_{h=1}^{N_{h}}\in\mathbb{R}^{N\times\frac{C_{\mathcal{E}}}{N_{h}}}$, in order to enhance the expressiveness and diversity of results without incurring extra computational costs.
The multi-head similarity function ($\operatorname{S}$) and cross attention layer ($\operatorname{MCA}$) are defined in Equations 6 and 7:
$\operatorname{S}(q_{i}^{h},k_{j}^{h})=\operatorname{softmax}\left(\frac{q_{i}^{h}{k_{j}^{h}}^{T}}{\sqrt{C_{\mathcal{E}}/N_{h}}},1\right),$ (6)
$\operatorname{MCA}(Q_{i},K_{j},V_{j})=\operatorname{Linear}(\operatorname{Concat}(\{\operatorname{S}(q_{i}^{h},k_{j}^{h})v_{j}^{h}\}_{h=1}^{N_{h}})),$ (7)
where $\operatorname{softmax}(\cdot,k)$ applies softmax operation across the $k$-th axis; $\operatorname{Linear}$ and $\operatorname{Concat}$ mean linear projection and channel-wise concatenation, respectively. Our method targets equations 6 and 7 for intuitive and efficient plane-specific modifications, achieving implicit plane matching and direct pose regression.
Plane Correspondence Probability Function.  To overcome the limitations of the multi-stage two-view plane reconstruction paradigm [7, 8, 9], here we specifically devise an inter-frame correspondence attention structure to learn reliable plane embeddings, which enables the network to autonomously acquire probabilities of plane correspondence and conduct pose inference in a single forward pass. This design also eliminates the dependency on either ground truth correspondence supervision or initial pose.
@@
$\small\operatorname{C}(Q_{i},K_{j})=\operatorname{softmax}(\frac{Q_{i}K_{j}^{T}}{\sqrt{C_{\mathcal{E}}}},1)\odot\operatorname{softmax}(\frac{Q_{i}K_{j}^{T}}{\sqrt{C_{\mathcal{E}}}},2){\color[rgb]{0,0,0},}$ (8)
We compute a 2D correspondence matrix $\operatorname{C}(Q_{i},K_{j})$, where the element at the $m^{th}$ row and $n^{th}$ column $\operatorname{C}_{mn}(Q_{i},K_{j})$ denotes the probability that the $m^{th}$ plane embedding from the $i^{th}$ image $I_{i}$ corresponds to the $n^{th}$ plane embedding from $j^{th}$ image $I_{j}$, indicating their likelihood of representing the same plane instance. For the task of two-view plane reconstruction, we stick to configurations of $\{i=1,j=2\}$ and $\{i=2,j=1\}$.
@@
Inter-frame Plane Aware Cross Attention.  The standard cross attention offers an efficient approach to selectively utilize one of the input plane sequences, but overlooks the interaction between two inputs. Therefore, we adopt bilinear attention [45, 25] to incorporate planar information from both views through the plane correspondence probability distribution $\operatorname{C}(Q_{i},K_{j})$:
$\operatorname{PCA}(Q_{i},K_{j},V_{i},V_{j})=$ $\operatorname{Linear}(\operatorname{Concat}($
$\{(v_{i}^{h})^{T}\operatorname{C}(Q_{i},K_{j})v_{j}^{h}\}_{h=1}^{N_{h}})){\color[rgb]{0,0,0},}$ (9)
Unlike the query and key terms, the value is still divided into $N_{h}$ segments along the feature dimension, maintaining the advantages of multi-head attention. The value segments share the correspondence attention (Equation 8) from unsplit query and key, allowing our model to selectively attend to actual corresponding plane embedding pairs across distinct sub-spaces.
As shown in Figure 4, two parallel plane aware cross attention layers (Equation 5.1) are used to capture integrated features of the corresponding planes from $I_{1}$ to $I_{2}$ as well as from $I_{2}$ to $I_{1}$. It is worth noting that pose ViT [25] employs the identical features on both sides of bilinear attention matrix. However, we intuitively choose to cross-place embedding sequences of distinct images to model correspondences, thereby enhancing learning efficiency and yielding improved results. We will validate various design disparities through subsequent ablation studies in Section 7.6.